OpenStudy (anonymous):

(2a^9-1/2b^9)^2 Multiply

OpenStudy (turingtest):

you mean\[\left({2a^9-1\over b^9}\right)\cdot\left({2a^9-1\over b^9}\right)\]??

OpenStudy (anonymous):